# DehazeGS: Seeing Through Fog with 3D Gaussian Splatting

### v4

@@ Abstract
Current novel view synthesis tasks primarily rely on high-quality and clear images. However, in foggy scenes, scattering and attenuation can significantly degrade the reconstruction and rendering quality. Although NeRF-based dehazing reconstruction algorithms have been developed, their use of deep fully connected neural networks and per-ray sampling strategies leads to high computational costs. Moreover, NeRF’s implicit representation struggles to recover fine details from hazy scenes. In contrast, recent advancements in 3D Gaussian Splatting achieve high-quality 3D scene reconstruction by explicitly modeling point clouds into 3D Gaussians. In this paper, we propose leveraging the explicit Gaussian representation to explain the foggy image formation process through a physically accurate forward rendering process. We introduce DehazeGS, a method capable of decomposing and rendering a fog-free background from participating media using only muti-view foggy images as input. We model the transmission within each Gaussian distribution to simulate the formation of fog. During this process, we jointly learn the atmospheric light and scattering coefficient while optimizing the Gaussian representation of the hazy scene. In the inference stage, we eliminate the effects of scattering and attenuation on the Gaussians and directly project them onto a 2D plane to obtain a clear view. Experiments on both synthetic and real-world foggy datasets demonstrate that DehazeGS achieves state-of-the-art performance in terms of both rendering quality and computational efficiency. https://dehazegs.github.io/
1 Introduction
@@
In recent years, Neural Radiance Fields (NeRF)  Mildenhall et al. (2021) have leveraged deep fully connected neural networks to implicitly represent 3D scenes and achieve impressive rendering quality through differentiable volumetric rendering techniques Levoy (1990); Max (1995). Subsequent works have focused on enhancing NeRF’s performance, particularly in terms of training speed  Chen et al. (2022); Müller et al. (2022) and rendering efficiency Garbin et al. (2021); Yu et al. (2021). A recent work, 3D Gaussian Splatting (3DGS)  Kerbl et al. (2023), abandons NeRF’s implicit rendering approach and instead represents scenes explicitly by converting point clouds into 3D Gaussians (ellipsoids). Thanks to its tile-based multi-threaded parallel rendering mechanism, 3DGS achieves high-quality real-time rendering.
@@
Figure 1: DehazeGS can generate accurate rendering results for scenes with scattering media $(b)$. By learning disentangled representations of the participating media and the clear scene, it simultaneously recovers the clear scene$(c)$ and obtains accurate depth estimations $(d)$.
@@
However, when capturing scenes containing scattering media, the light received by the detector is disrupted by the medium. The detected light mainly consists of two components: first, light reflected from the surfaces of objects that has been attenuated by particles of medium, and second, ambient light (e.g., sunlight) scattered by the particles of medium. This interference often leads to scenes suffering from low contrast and limited visibility. 3DGSs and NeRFs, most methods are designed for clear media, which is disadvantageous for applications such as autonomous driving and robotic operations under foggy scene conditions. Although some dehazing reconstruction methods based on NeRF  Ramazzina et al. (2023); Li et al. (2023) have been proposed, these approaches either suffer from extremely slow training and rendering speeds or are limited to specific indoor scenes. In contrast, traditional dehazing algorithms in low-level computer vision Zhang and Patel (2018); Yang et al. (2022), typically require large-scale paired or unpaired datasets for training. Moreover, whether single-image or multi-image dehazing, these methods fail to meet the requirements of multi-view consistency necessary for 3D reconstruction. Most methods remain confined to 2d image plane Dong et al. (2020); Zheng et al. (2021); Qin et al. (2020), lacking consideration of 3D spatial information. Currently, there are no effective methods, and the ability to perform physically accurate modeling and to separate scattering in participating media is crucial for imaging and scene understanding tasks.
@@
To address the above challenges, we propose DehazeGS, the first differentiable physics-based 3DGS dehazing model. We have pioneered a novel representation of 3DGS for foggy scenes. We conceive the idea of modeling Gaussians as scattering particles in the atmosphere. Our method is primarily divided into three components. 1) We establish a mapping relationship between the Gaussian ellipsoid depth $G(z)$ and the transmission Gaussian $G(t)$. By querying the depth of Gaussians as input to a convolutional neural network, we obtain the transmission Gaussian. The scattering coefficients are stored as the weights of the convolutional neural network. 2) After obtaining the transmission Gaussian, we model the foggy Gaussians using the Atmospheric Scattering Model (ASM) McCartney (1976); Narasimhan and Nayar (2002), in conjunction with the estimated atmospheric light coefficients. Once the foggy Gaussians are rasterized, the estimated fog map can be rendered. Using the real fog map as supervision, we jointly optimize the parameters of the Gaussians while training the neural network. 3) As the scene depth increases, light attenuation becomes more pronounced. The accuracy of depth is crucial for estimating the transmission map, and the accuracy of the transmission map, in turn, impacts the dehazing performance. Therefore, to obtain more precise depth estimation, we apply depth regularization by using pseudo-depth maps (as shown in Figure 2) as priors when optimizing each input image. In addition, to emphasize the restoration of distant details, we incorporated a depth-weighted reconstruction loss. Note that for the transmission Gaussian obtained in 1), we directly perform alpha blending to render the transmission map. We use the transmission map generated by the dark channel prior  He et al. (2010) and the bright channel prior Zhang et al. (2021) as supervision to guide the optimization of the rendered transmission map.
Our contributions can be summarized as follows:
1) We propose the first framework for learning clear 3D Gaussian splatting solely from multi-view foggy images. This framework is capable of learning disentangled representations of participating media and clear scenes, achieving fast, high-quality dehazing reconstruction.
2) We propose a novel representation, where atmospheric scattering and clear content are modeled separately. A transmission is established on each Gaussian distribution. Based on the ASM model, the learnable global atmospheric light is combined with the transmission Gaussian and further applied to the latent clear Gaussian, making the Gaussians represent foggy scenes.
3) A novel prior loss committee is introduced, incorporating relevant prior information to effectively enhance the model’s dehazing performance and achieve more accurate depth estimation.
4) Our method achieves real-time dehazing rendering, with most scenes requiring only 3,000 iterations to reach optimal results, and the training time taking approximately 1 minute. Compared to NeRF-based methods, our approach reduces the time overhead by several hundred times.
2 Related Works
2.1 Dehazing Based on Image Processing
In traditional image processing, there are numerous dehazing algorithms that typically leverage the statistical properties of clear images to estimate the transmission map and atmospheric light. These estimates are then used in the inverse process of the ASM to recover the underlying clear image. One of the most classic algorithms is the Dark Channel Prior (DCP) He et al. (2010) dehazing algorithm, which is based on statistical observations from a large number of outdoor haze-free scenes, the majority of local patches contain some pixels with very low intensity in at least one color channel. Subsequently, many other related algorithms have also been developed, such as the Bright Channel Prior Zhang et al. (2021) and the Color Attenuation Prior Zhu et al. (2015). With the advancement of deep learning, neural networks have gradually been applied to image processing, which can be categorized into physics-based approaches and end-to-end methods. In physics-based algorithms, neural networks are primarily used to estimate the atmospheric light and transmission map of hazy images. For example, DCPDN Zhang and Patel (2018) proposed a densely connected pyramid network capable of jointly estimating atmospheric light and transmission maps. PSD Chen et al. (2021) introduced a dehazing network guided by physical priors. For end-to-end dehazing methods, for example, DehazeNet Cai et al. (2016) utilizes a CNN to learn the mapping relationship between the original hazy images and the corresponding medium transmission maps. Dong et al. (2020) proposed a multi-scale enhanced dehazing network with dense feature fusion, which is based on a U-net architecture and directly takes hazy images as input, processing them through multiple network layers to produce clear images.
2.2 Neural Radiance Fields for Dehazing
NeRF Mildenhall et al. (2021) uses a fully connected neural network to represent a scene, where the input is the 5D coordinates of the sampled points, and the output is the volumetric density and view-dependent radiance (color) at those spatial locations. Finally, the color and density of the sampled points are integrated along the view direction to produce colors on the image plane using volumetric rendering techniques. ScatterNeRF Ramazzina et al. (2023) builds upon NeRF by introducing an additional MLP to learn the opacity of the scattering medium and its associated color. A scattering term is added to the volumetric rendering equation, blending clear and blurred sampled points to render foggy images. Its primary limitation lies in the need for high-density sampling along each ray in space, which requires processing through a deep fully connected network to compute the volumetric density and color of sampled points. DehazeNeRF Chen et al. (2023) extends the volumetric rendering equation by simulating the physical reality of atmospheric scattering. It also combines several regularization strategies to achieve 3D shape reconstruction while removing haze. However, this method is primarily designed for indoor scenes. The network in Dehazing-NeRF Li et al. (2023) that estimates the scattering coefficient and atmospheric light requires pre-training, and its framework cannot achieve joint optimization of scattering and clear content. Additionally, it is limited to synthetic datasets, each containing a single object (e.g., lego and chair).
3 Method
3.1 Atmospheric Scattering Model
@@
The Atmospheric Scattering Model (ASM) McCartney (1976); Narasimhan and Nayar (2002), is a physical model that describes the changes in light as it propagates through the atmosphere due to scattering and absorption. Based on the principles of light propagation, it reveals how particles in the atmosphere affect the quality of captured images. Its mathematical formulation is expressed as:
$I{(x)}=J(x)t(x)+A(1-t(x)).$ (1)
@@
Here, $I$ represents the observed hazy image, $J$ denotes the underlying clear image, and $A$ is the global atmospheric light, which is generated by the scattering of ambient light (typically sunlight) by participating media, represents the background light intensity received by imaging devices. It primarily affects the overall brightness of the image. $t(x)$ represents the transmission map, where $x$ denotes the position of a pixel. The transmission map primarily serves to characterize the attenuation of light as it travels through scattering media (fog) in the atmosphere, directly influencing the accuracy of the dehazing results. $t(x)$ can be further expressed as $t(x)=e^{-\beta d(x)}$, where $\beta$ represents the scattering coefficient and $d(x)$ denotes the scene depth at pixel $x$.
3.2 3D Gaussian Splatting
@@
The 3D Gaussian approach does not rely on neural radiance fields. instead, it represents the scene as a series of 3D Gaussian distributions Yu et al. (2024). Based on the initialized sparse point cloud, a set of 3D Gaussians, defined as $G$, is parameterized by its 3D coordinates $\boldsymbol{x}\in\mathbb{R}^{3}$, 3D covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $\alpha\in\mathbb{R}$ and color $c\in\mathbb{R}^{3}$. $c$ is represented by spherical harmonics for view-dependent appearance. The distribution of each Gaussian is defined as:
@@
3D gaussians are projected into the 2D image space from a given camera pose $\boldsymbol{P}_{\boldsymbol{c}}=\{\boldsymbol{R}_{\boldsymbol{c}}\in\mathbb{R}^{3\times 3},\boldsymbol{~{}}\boldsymbol{t}_{\boldsymbol{c}}\in\mathbb{R}^{3}\}$. Given the viewing transformation $\boldsymbol{W}$ and 3D covariance matrix $\boldsymbol{\Sigma}$, the projected 2D covariance matrix $\boldsymbol{\Sigma}^{\prime}$ is computed using, as described in Zwicker et al. (2001):
@@
Subsequently, the transformed Gaussians are sorted based on their depth and the sorted Gaussians are rasterized to render pixel values using the following volume rendering equation:
@@
where $c_{i}$ is the learned color and the $\alpha^{{}^{\prime}}_{i}$ is the multiplication result of the learned opacity $\alpha_{i}$ and the 2D Gaussian.
3.3 DehazeGS
Modeling Gaussian distributions in foggy scenes. Fog is a natural phenomenon caused by the scattering effect of aerosol particles in the atmosphere. When photographing in foggy conditions, the light received by the sensor primarily originates from two sources: the light reflected from the object’s surface, which undergoes scattering and attenuation by the particles, along with the environmental light (such as sunlight), which is also scattered by the particles. As light passes through the scattering medium, it experiences varying degrees of scattering and attenuation, with the attenuation increasing as the scene depth increases (i.e., the transmission rate decreases). Since the scene represented by 3DGS is modeled using 3D Gaussians, we propose directly defining the transmission function $t(x)$ on each Gaussian distribution. The physical meaning of this function is the proportion of light intensity that remains unattenuated after passing through the particles. Furthermore, the transmission function is associated with the depth of the Gaussian. For any Gaussian $\mathcal{G}(\boldsymbol{x})$, its depth $d_{\mathcal{G}}(\boldsymbol{x})$ (represents the depth $d$ of the Gaussian ellipsoid at coordinate $x$) is queried, and its corresponding transmission Gaussian $t_{\mathcal{G}}(\boldsymbol{x})$ (represents the transmittance $t$ of the Gaussian ellipsoid at coordinate $x$) is obtained through a single-layer one-dimensional convolutional neural network. The mapping relationship can be expressed as $\left.F_{\beta}:{d_{\mathcal{G}}(\boldsymbol{x})}\rightarrow t_{\mathcal{G}}{(\boldsymbol{x})}\right.$, where $\beta$ represents the learnable weight, that is, the scattering coefficient. The detailed structure of the network is shown in the upper-left part of Fig. 2. Since the transmission range is between 0 and 1, with smaller values indicating more severe attenuation of light, we apply a sigmoid ($\sigma$) activation function at the output of the Gaussian transmission. To ensure the non-negativity of the product of depth and scattering coefficient, we apply a ReLU activation function to the exponential term. Finally, based on equation $t(x)=e^{-\beta d(x)}$ from Section 3.1, we can obtain the transmission function formula based on Gaussians as follows,
@@ 3.3 DehazeGS
This approach of directly establishing the transmission function and atmospheric light on the Gaussians effectively avoids the high sampling and rendering costs associated with NeRF-based method Chen et al. (2023); Ramazzina et al. (2023). Moreover, this approach effectively leverages the reconstruction capability in 3DGS, enhancing the structural consistency of the dehazed scene across multiple viewpoints.
After rendering the fog image, the reconstruction loss between the rendered image and the real fog image is calculated:
@@
Transmission Gaussian Optimization. We directly perform alpha blending on the transmission Gaussians $\mathcal{G}(\boldsymbol{x},t)$ to obtain the rendered transmittance map $\mathcal{\hat{T}}(\boldsymbol{P})$:
@@
where $\boldsymbol{P}$ represents pose. We propose to use the Dark Channel Prior (DCP) He et al. (2010) and further consider Bright Channel Prior (BCP) Sun and Guo (2016) algorithms to guide the optimization of our estimated transmission map. The DCP is one of the most well-known dehazing algorithms, leveraging the statistical properties of clear images to estimate the transmission map and atmospheric light. we follow the prior of traditional methods Chen et al. (2021); Golts et al. (2019) and reformulate the prior as a 3DGS-based energy function:
$\begin{split}\mathcal{L}_{DCP}&={\mathcal{{T}}}^{T}(\boldsymbol{P})L\mathcal{{T}}(\boldsymbol{P})+\lambda{\tilde{\mathcal{T}}(\boldsymbol{P})}^{T}\tilde{\mathcal{T}}(\boldsymbol{P}),\\ \tilde{\mathcal{T}}(\boldsymbol{P})&=(\mathcal{T}(\boldsymbol{P})-\mathcal{\hat{T}}{(\boldsymbol{P})}).\\ \end{split}$ (10)
Here, $\mathcal{T}(\boldsymbol{P})$ represents the transmission map under the corresponding pose $\boldsymbol{P}$, estimated by the DCP algorithm. To mitigate the darkening effect caused by DCP, we further introduce BCP, a brightness enhancement algorithm based on the bright channel prior, which helps to enhance the overall brightness of the rendered dehazed images. The loss function is formulated as follows,
$\left.\mathcal{L}_{BCP}=~{}\middle|\middle|\mathcal{T}(\boldsymbol{P})\left.-\mathcal{\hat{T}}{(\boldsymbol{P})}\middle|\right|_{1}\right.$ (11)
where $\mathcal{T}(\boldsymbol{P})$ is estimated from the BCP.
@@
To guide the optimization of Gaussian ellipsoid depths, we utilize DepthAnything V2 Yang et al. (2024b) to predict the depth of foggy images, generating a pseudo-depth map to guide the optimization of the 3DGS-rendered depth. We found that, in a very small number of scenarios, introducing the Contrast Limited Adaptive Histogram Equalization (CLAHE) can slightly improve performance. However, in most cases, CLAHE is not required, and directly estimating the depth from multi-view foggy images as supervision is sufficient.
@@
Furthermore, to enhance the recovery of distant details, we incorporate a depth-weighted reconstruction loss $\mathcal{L}_{\hat{D}_{recon}}$:
@@
4 Experiments
4.1 Experimental Setup
@@
Synthetic Foggy Dataset. We selected four representative scenes from the Mip-NeRF Barron et al. (2021) dataset, including two indoor scenes (bonsai and counter) and two outdoor scenes (garden and stump). For each scene, a small subset of images was reserved as the test set, while the remaining images were processed as follows: we first employed a depth estimation network Yang et al. (2024b) to estimate the ground truth depth of the clear images, then randomly generated scattering coefficients and global atmospheric light for each scene. Finally, the images were fogged based on the principles of the atmospheric scattering model.
@@
Real Foggy Dataset. We used three indoor foggy scenes provided by DehzeNeRF Chen et al. (2023), captured with two professional fog machines and one mobile phone. The three scenes are bear, elephant, and lion. Each scene contains 82, 58, and 68 foggy images and 47, 79, and 47 clear images, respectively. We combined the clear and foggy images from the same scene and input them into COLMAP to obtain the sparse point cloud and corresponding poses.

### v5

@@ Abstract
Current novel view synthesis methods are typically designed for high-quality and clean input images. However, in foggy scenes, scattering and attenuation can significantly degrade the quality of rendering. Although NeRF-based dehazing approaches have been developed, their reliance on deep fully connected neural networks and per-ray sampling strategies leads to high computational costs. Furthermore, NeRF’s implicit representation limits its ability to recover fine-grained details from hazy scenes. To overcome these limitations, we propose learning an explicit Gaussian representation to explain the formation mechanism of foggy images through a physically forward rendering process. Our method, DehazeGS, reconstructs and renders fog-free scenes using only multi-view foggy images as input. Specifically, based on the atmospheric scattering model, we simulate the formation of fog by establishing the transmission function directly onto Gaussian primitives via depth-to-transmission mapping. During training, we jointly learn the atmospheric light and scattering coefficients while optimizing the Gaussian representation of foggy scenes. At inference time, we remove the effects of scattering and attenuation in Gaussian distributions and directly render the scene to obtain dehazed views. Experiments on both real-world and synthetic foggy datasets demonstrate that DehazeGS achieves state-of-the-art performance.
Introduction
@@
In recent years, Neural Radiance Fields (NeRF)  (Mildenhall et al. 2021) have leveraged deep fully connected neural networks to implicitly represent 3D scenes and achieve impressive rendering quality through differentiable volumetric rendering techniques (Levoy 1990; Max 1995). Subsequent works have focused on enhancing NeRF’s performance, particularly in terms of training speed  (Chen et al. 2022; Müller et al. 2022) and rendering efficiency (Garbin et al. 2021; Yu et al. 2021). A recent work, 3D Gaussian Splatting (3DGS)  (Kerbl et al. 2023), abandons NeRF’s implicit rendering approach and instead represents scenes explicitly by converting point clouds into 3D Gaussians (ellipsoids). Thanks to its tile-based multi-threaded parallel rendering mechanism, 3DGS achieves high-quality real-time rendering.
@@
Figure 1: DehazeGS can generate accurate rendering results for scenes with participating media $(b)$. By learning disentangled representations for the participating media and the underlying scene, it simultaneously recovers the clear scene $(c)$ and obtains accurate depth estimation $(d)$.
@@
However, when capturing scenes containing scattering media, the light received by the detector is disrupted by the medium. The detected light mainly consists of two components: first, light reflected from the surfaces of objects that has been attenuated by particles of medium, and second, ambient light scattered by the particles of medium. This interference often leads to scenes suffering from low contrast and limited visibility. For both 3DGS and NeRF, most methods are designed for clear media, which is disadvantageous for applications such as autonomous driving and robotic operations under foggy scene conditions. Although some dehazing reconstruction methods based on NeRF  (Ramazzina et al. 2023; Li et al. 2023) have been proposed, these approaches either suffer from extremely slow training and rendering speeds or are limited to specific indoor scenes. In contrast, traditional dehazing algorithms in low-level computer vision (Zhang and Patel 2018; Yang et al. 2022; Zhang, Zhou, and Li 2024) typically require large-scale paired or unpaired datasets for training. Moreover, these methods remain confined to a 2D image plane (Dong et al. 2020; Zheng et al. 2021; Qin et al. 2020), lacking consideration of 3D spatial information. Therefore, they fail to meet the requirements of multi-view consistency across viewpoints necessary for 3D reconstruction. Feeding such inconsistent images into reconstruction models significantly degrades reconstruction quality (Detailed proofs can be found in the supplementary materials or Fig. 3). Thus, in the field of novel view synthesis, there is an urgent need for an end-to-end dehazing reconstruction model capable of effectively achieving physically accurate modeling and separating scattering effects in participating media.
To address the above challenges, we propose DehazeGS, the first differentiable physics-based 3DGS dehazing model. We have pioneered a novel representation of 3DGS for foggy scenes. Our method is primarily divided into three components. 1) Inspired by the transmission formula $t(x)=e^{-\beta d(x)}$ of the Atmospheric Scattering Model (ASM) (McCartney 1976; Narasimhan and Nayar 2002), we propose to establish a mapping relationship between Gaussian depth $d_{\mathcal{G}}(\boldsymbol{x})$ and Gaussian transmission $t_{\mathcal{G}}(\boldsymbol{x})$. By querying the depth of Gaussians as input to a convolutional neural network, we obtain the Gaussian transmission. The scattering coefficients are stored as the weights of the convolutional neural network. 2) After obtaining the Gaussian transmission, we model the foggy Gaussians based on the Atmospheric Scattering Model (ASM), in conjunction with the estimated atmospheric light coefficients. Once the foggy Gaussians are rasterized, the estimated fog map can be rendered. Using the real foggy map as supervision, we jointly optimize the parameters of the Gaussians while training the neural network and fully render the foggy scenes. 3) The accuracy of depth is crucial for estimating the transmission. Therefore, while optimizing the Gaussian representation, we adopt pseudo-depth maps as priors to impose depth regularization. To emphasize the restoration of distant details, we incorporate a depth-weighted reconstruction loss. In addition, note that for the Gaussian transmission obtained in the first stage, we directly perform alpha blending to render the transmission map. We propose to use the transmission map generated by the dark channel prior (He, Sun, and Tang 2010) and the bright channel prior (Zhang, Wang, and Wang 2021) as supervision to guide the optimization of the rendered transmission map.
Our main contributions are as follows:
  • We propose the first framework for learning clear 3D Gaussian splatting solely from multi-view foggy images. This framework is capable of learning disentangled representations of participating media and clear scenes, achieving fast, high-quality dehazing reconstruction.
  • We establish a mapping between Gaussian depth and its transmission, and based on the Atmospheric Scattering Model (ASM), we propose a representation method for Gaussian primitives in foggy scenes.
  • We propose a novel Gaussian-based prior loss committee, leveraging these physical priors to guide the fine-tuning of our model. The combination of these prior methods further enhances the dehazing rendering performance of our model.
  • Experiments on synthetic and real foggy datasets demonstrate that our method surpasses existing approaches in both rendering quality and speed, achieving state-of-the-art results.
Related Works
Dehazing Based on Image Processing
In the field of image processing, one of the most classic dehazing algorithms is the Dark Channel Prior (DCP) algorithm (He, Sun, and Tang 2010), which is based on statistical observations from a large number of outdoor haze-free scenes. Subsequently, many other related algorithms have also been developed, such as the Bright Channel Prior (Zhang, Wang, and Wang 2021) and the Color Attenuation Prior (Zhu, Mai, and Shao 2015). With deep learning advancement, neural networks are increasingly applied to image processing, categorized into physics-based approaches and end-to-end methods. In physics-based algorithms, they primarily estimate hazy images’ atmospheric light and transmission maps. For example, DCPDN (Zhang and Patel 2018) proposed a densely connected pyramid network capable of jointly estimating atmospheric light and transmission maps. PSD (Chen et al. 2021) introduced a dehazing network guided by physical priors. For end-to-end dehazing methods, for example, DehazeNet (Cai et al. 2016) utilizes a CNN to learn the mapping relationship between the original hazy images and the corresponding medium transmission maps. AOD-Net (Li et al. 2017) is a lightweight end-to-end dehazing network that can directly generate haze-free images without estimating transmission maps or atmospheric light independently. MSBDN (Dong et al. 2020) proposed a multi-scale enhanced dehazing network with dense feature fusion. DehazeFormer(Song et al. 2023) is designed as an end-to-end single-image dehazing method based on the Swin Transformer architecture.
Neural Radiance Fields for Dehazing
NeRF (Mildenhall et al. 2021) employs a deep fully connected neural network for scene representation, which takes 5D sampled point coordinates as input and outputs volumetric density and view-dependent radiance (i.e., color). Volumetric rendering is then used to integrate the color and density along each viewing ray to generate pixel colors. ScatterNeRF (Ramazzina et al. 2023) builds upon NeRF by introducing an additional MLP to learn the opacity of the scattering medium and its associated color. A scattering term is added to the volumetric rendering equation, blending clear and blurred sampled points to render foggy images. Its primary limitation lies in the need for high-density sampling along each ray in space, which requires processing through a deep fully connected network to compute the volumetric density and color of sampled points. DehazeNeRF (Chen et al. 2023) extends the volumetric rendering equation via simulating atmospheric scattering physics, integrating regularization strategies for dehazing and 3D shape reconstruction. However, it is mainly designed for indoor scenes. Dehazing-NeRF’s (Li et al. 2023) scattering coefficient/atmospheric light estimation network needs pre-training, lacking joint optimization of scattering medium and clear content. Additionally, it is confined to synthetic datasets with single objects (e.g., Lego, chair).
@@
Method
Preliminary
@@
Atmospheric Scattering Model. The Atmospheric Scattering Model (McCartney 1976; Narasimhan and Nayar 2002) (ASM) is a physical model that describes the changes in light as it propagates through the atmosphere due to scattering and absorption. Based on the principles of light propagation, it reveals how particles in the atmosphere affect the quality of captured images. Its mathematical formulation is expressed as:
I(x)=J(x)t(x)+A(1t(x)).I{(x)}=J(x)t(x)+A(1-t(x)). (1)
@@
Here, $I$ represents the observed hazy image, $J$ denotes the underlying clear image, and $A$ is the global atmospheric light, which represents the background light intensity received by the imaging devices. It is generated by the scattering of ambient light within participating media and primarily affects the overall brightness of the image. $t(x)$ represents the transmission map, where $x$ denotes the position of a pixel. The transmission map primarily serves to characterize the attenuation of light as it travels through scattering media (fog) in the atmosphere, directly influencing the accuracy of the dehazing results. $t(x)$ can be further expressed as $t(x)=e^{-\beta d(x)}$, where $\beta$ represents the scattering coefficient and $d(x)$ denotes the scene depth at pixel $x$.
@@
3D Gaussian Splatting. The 3D Gaussian approach does not rely on neural radiance fields. Instead, it represents the scene as a series of 3D Gaussian distributions (Yu et al. 2024). Based on the initialized sparse point cloud, a set of 3D Gaussians, defined as $G$, is parameterized by its 3D coordinates $\boldsymbol{x}\in\mathbb{R}^{3}$, 3D covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $\alpha\in\mathbb{R}$ and color $\in\mathbb{R}^{3}. cc$ is represented by spherical harmonics for view-dependent appearance. The distribution of each Gaussian is defined as:
@@
3D gaussians are projected into the 2D image space from a given camera pose $\boldsymbol{P}_{\boldsymbol{c}}=\{\boldsymbol{R}_{\boldsymbol{c}}\in\mathbb{R}^{3\times 3},\boldsymbol{~}\boldsymbol{t}_{\boldsymbol{c}}\in\mathbb{R}^{3}\}$. Given the viewing transformation $\boldsymbol{W}$ and 3D covariance matrix $\boldsymbol{\Sigma}$, the projected 2D covariance matrix $\boldsymbol{\Sigma}^{\prime}$ is computed using, as described in (Zwicker et al. 2001):
@@
Subsequently, the transformed Gaussians are sorted based on their depth, and the sorted Gaussians are rasterized to render pixel values using the volume rendering equation:
@@
where $c_{i}$ is the learned color and the $\alpha^{\prime}_{i}$ is the multiplication result of the learned opacity $\alpha_{i}$ and the 2D projected Gaussion.
DehazeGS
Modeling Gaussian distributions in foggy scenes. Fog is a natural phenomenon caused by the scattering effect of aerosol particles in the atmosphere. When photographing in foggy conditions, the light received by the sensor primarily originates from two sources: the light reflected from the object’s surface, which undergoes scattering and attenuation by the particles, along with the environmental light, which is also scattered by the particles. As light passes through the scattering medium, it experiences varying degrees of scattering and attenuation, with the attenuation increasing as the scene depth increases (i.e., the transmission rate decreases). Since the scene represented by 3DGS is modeled using 3D Gaussians, we propose directly defining the transmission function $t(x)$ on each Gaussian distribution. The physical meaning of this function is the proportion of light intensity that remains unattenuated after passing through the particles. Furthermore, the transmission function is associated with the depth of the Gaussian. For any Gaussian $\mathcal{G}(\boldsymbol{x})$, its depth $d_{\mathcal{G}}(\boldsymbol{x})$ (represents the depth $d$ of the Gaussian ellipsoid at coordinate $x$) is queried, and its corresponding transmission $t_{\mathcal{G}}(\boldsymbol{x})$ (represents the transmittance $t$ of the Gaussian ellipsoid at coordinate $x$) is obtained through a single-layer one-dimensional convolutional neural network. The mapping relationship can be expressed as $\left.F_{\beta}:{d_{\mathcal{G}}(\boldsymbol{x})}\rightarrow t_{\mathcal{G}}{(\boldsymbol{x})}\right$., where $\beta$ represents the learnable weight, that is, the scattering coefficient. The detailed structure of the network is shown in the upper-left part of Fig. 2.
Since the $t_{\mathcal{G}}(\boldsymbol{x})$ ranges between 0 and 1, with smaller values indicating more severe light attenuation, we apply a ReLU activation function to the exponential term to ensure the non-negativity of $\beta d_{\mathcal{G}}(\boldsymbol{x})$ (the product of the scattering coefficient $\beta$ and gaussian depth), thereby restricting the output range of the function to between 0 and 1. However, in practical experiment, we found that when ensuring the non-negativity of $\beta d_{\mathcal{G}}(\boldsymbol{x})$, feeding the result of the entire negative exponential function into a sigmoid activation function can improve the training stability of the model. Although adding a sigmoid is not strictly necessary.
Finally, based on equation $t(x)=e^{-\beta d(x)}$ from Section 3.1, we can obtain the transmission function formula based on Gaussians as follows:
@@
This approach of directly establishing the transmission function and atmospheric light on the Gaussians effectively avoids the high sampling and rendering costs associated with NeRF-based methods (Chen et al. 2023; Ramazzina et al. 2023). Moreover, this approach effectively leverages the reconstruction capability in 3DGS, enhancing the structural consistency of the dehazed scene across multiple viewpoints.
After rendering the fog image, the reconstruction loss between the rendered image $\hat{I}$ and the real fog image $I$ is calculated. Consistent with the original 3D Gaussian reconstruction loss, The loss function is $\mathcal{L}_{1}$ combined with a DSSIMD-SSIM term, $\lambda$ = 0.2:
@@
Gaussian Transmission Optimization. We directly perform alpha blending on the transmission values $t_{\mathcal{G}}(\boldsymbol{x})$ of each Gaussian to obtain the rendered transmittance map $\mathcal{\hat{T}}(\boldsymbol{P})$:
@@
where $\boldsymbol{P}$ represents pose. We propose to use the Dark Channel Prior (DCP) (He, Sun, and Tang 2010) and further consider Bright Channel Prior (BCP) (Sun and Guo 2016) algorithms to guide the optimization of our estimated transmission map. The DCP is one of the most well-known dehazing algorithms, leveraging the statistical properties of clear images to estimate the transmission map and atmospheric light. we follow the prior of traditional methods (Chen et al. 2021; Golts, Freedman, and Elad 2019) and reformulate the prior as a 3DGS-based energy function:
[…]
Here, $\mathcal{T}_{D}(\boldsymbol{P})$ represents the transmission map under the corresponding pose $\boldsymbol{P}$, estimated by the DCP algorithm. $L$ and $\lambda$ are the Laplacian-like matrix and hyper-parameter, respectively.
To mitigate the darkening effect caused by DCP, we further introduce BCP, a brightness enhancement algorithm based on the bright channel prior, which helps to enhance the overall brightness of the rendered dehazed images. The loss function is formulated as follows,
$\left.\mathcal{L}_{BCP}=~\middle|\middle|\mathcal{T}_{B}(\boldsymbol{P})\left.-\mathcal{\hat{T}}{(\boldsymbol{P})}\middle|\right|_{1}\right$. (11)
where $\mathcal{T}_{B}(\boldsymbol{P})$ is estimated from the BCP.
@@
We leverage DepthAnything V2 (Yang et al. 2024) to predict the depth maps from foggy images, generating pseudo-depth maps to guide the optimization of the 3DGS-rendered depth. Since the pseudo-depth maps are estimated from the foggy inputs and inherently contain certain errors, the alignment between $D_{pseudo}$ and $\hat{D}$ is relative rather than absolute. To address this, we apply weight decay during training to gradually reduce the weight of this loss. Specifically, we adopt a continuous learning rate decay schedule. The initial weight is set to 1 at step 0 and gradually reduced to 0.01 at the specified maximum step, with logarithmic interpolation applied between these steps.
@@
Furthermore, to enhance the recovery of distant details, we incorporate a depth-weighted reconstruction loss $\mathcal{L}_{d_{rec}}$:
@@
Experiments
Experimental Setup
@@
Synthetic Foggy Dataset. We chose four representative scenes from the Mip-NeRF (Barron et al. 2021) dataset, including two indoor scenes (bonsai and counter) and two outdoor scenes (garden and stump). For each scene, we employed the Depth Anything V2 model (Yang et al. 2024) to estimate ground-truth depths of original clear images. Subsequently, we stochastically sampled scattering coefficients and global atmospheric light for each scene. Finally, haze synthesis was performed on the images based on principles of the atmospheric scattering model. Additionally, we performed dehazing tests on a driving scene from the Waymo (Sun et al. 2020) dataset.
@@
Real Foggy Dataset. We used three indoor foggy scenes provided by DehzeNeRF (Chen et al. 2023), captured with two professional fog machines and one mobile phone. The three scenes are bear, elephant, and lion. Each scene contains a certain number of clear views and foggy views. We combined the clear and foggy views from the same scene and input them into COLMAP to obtain the sparse point cloud and corresponding poses.
Evaluation methods. During the training process, we only take the multi-view foggy views of the scene as input to render clear views, and compute evaluation metrics (PSNR / SSIM / LPIPS) using the rendered clear images and their corresponding real clear images.